Bob wants to have concrete driveway in fron of his house. The driveway will measure 58 2/3 feet long and 18 1/2 feet wide; and will be 3/4 foot thick. How many cubic yards of cement will be needed for the driveway? Round your answer to the nearest cubic yard.
Answer by Alan3354(21583)

Bob wants to have concrete driveway in fron of his house. The driveway will measure 58 2/3 feet long and 18 1/2 feet wide; and will be 3/4 foot thick. How many cubic yards of cement will be needed for the driveway? Round your answer to the nearest cubic yard.
---------------
58 2/3 ft = 704 inches
18 1/2 ft = 222 inches
3/4 foot = 9 inches
V = 704*222*9 cubic inches
V = 1406592 in^3
V = 1406592/1728 ft^3
V = 814 cubic ft
V = 814/27 cubic yr
V = ~ 30 cubic yards
-------------------
Suppose the problem was in metric units.
1.788 meters
5.63 meters
0.2286 meters
V = 17.88*5.63*0.2286 cubic meters
V = 23 m^3
Much simpler.